# A non-linear quantum neural network framework for entanglement engineering

### v1

@@ IV.1 NISQ scenario
A.M.P. acknowledges support by MUR (Ministero dell’Università e della Ricerca) through the PNRR Project ICON-Q-Partenariato EstesoNQSTI - PE00000023 - Spoke 2 - CUP: J13C22000680006.
@@
In the last years, a revival of photonic applications to quantum state properties (or features) extraction is showing a revival, with several works on photonic platforms [ExperimentalPropertiesPhotonicPlatform, ReviewReservoir-Innocenti2023, Reservoir-Spagnolo2022].
Appendix A The beam splitter reflectivity
@@
Appendix B Structure of the Response Function
@@
$\omega=\frac{2\pi}{T_{\mathrm{osc}}}$
@@
$\frac{T_{\mathrm{osc}}}{T_{\mathrm{int}}\,4\pi}=\frac{1}{2\omega T_{\mathrm{int}}}$
@@
In systems exhibiting memory-dependent (or “memristive”) behavior, the present response does not depend solely on the instantaneous value of the driving field, but also on its history within a specific temporal window. Equation (12) captures exactly this effect: $R(t)$ measures the deviation introduced by integrating the oscillatory signal over an interval $T_{\mathrm{int}}$, effectively encoding a temporal “lag” between the free oscillation and its memory-processed counterpart.
@@
This formulation offers a compact yet powerful tool to quantify the interplay between coherence, periodic driving, and memory depth. In the context of photonic or quantum memristive systems, the response function (12) is particularly useful for characterizing the temporal retention of coherence and for distinguishing between purely dissipative and genuinely history-dependent dynamics.
Limiting Cases
Small integration window.
When $T_{\mathrm{int}}\ll T_{\mathrm{osc}}$, Taylor expansion yields
$R(t)\approx\frac{T_{\mathrm{int}}}{2T_{\mathrm{osc}}}\cos\!\left(\frac{2\pi t}{T_{\mathrm{osc}}}\right)+\mathcal{O}\!\left(\frac{T_{\mathrm{int}}^{3}}{T_{\mathrm{osc}}^{3}}\right),$
showing that the response becomes approximately proportional to the instantaneous cosine of the oscillation.
Integration equal to one period.
If $T_{\mathrm{int}}=T_{\mathrm{osc}}$, then the two sine terms cancel exactly, leading to $R(t)=0$. This reflects the fact that integrating a full period of a sinusoid yields zero net displacement.
High-frequency limit.
As $T_{\mathrm{osc}}\to 0$ (i.e., $\omega\to\infty$), rapid oscillations suppress the response, and
$R(t)\to 0,$
consistent with the averaging out of high-frequency excitations over the finite integration window.
@@
B.1 Pure states random topologies analysis
Figure 7: Frames (a) (b). Histogram of the loss function achieved by a total of $10^{5}$ random $\mathcal{T}$ for 5 qubits. Frames (c) (d), histogram for 60 for 11 qubits. We compare the $\mathcal{U}_{\rm sc+1}^{\rm RN}$ (red bars) against $\mathcal{U}^{\theta}_{\rm RN}$ (yellow bars). The network loss function maximizes the MW values of the output states by minimizing eq.6. So, values close to zero in the loss function indicate higher values of global entanglement in the output pure state. We can clearly see that the presence of a non-linearity as given by eq.13 response function, unlocks a greater optimal minima search.
As Fig.7 unfolds, the non-linear components unlock better outcomes, with many more random QNN that can achieve values of MW above 0.8. This fact can be leveraged to prepare a training dataset for classical deep learning solutions meant to synthesize optimal topologies. We also run a quick analysis for a small noise scenario, where only amplitude damping is considered. Again, we confirm that the presence of non-linear functions is necessary to achieve higher performances; the results are shown in Sec.D. Another natural question that arises is the dependency of the networks on initial conditions and parameter variation, i.e., the $T_{osc/int}$ that describes the physical gate, even though we are in a noiseless case. Results of the noiseless study can be found in C and show that for this scenario, no dependency is found.
The topologies of 0.8 threshold – As shown in panel (a), the MW value of 0.8 turns out as a threshold, or a specific boundary when a transition occurs. We analyze the structure of topologies below 0.8 and found that all of them have at least one wire set (A,B) with $A>B$ and that connect qubits that are not neighbors. This will play a critical role in Sec.IV.1 when the presence of noise will be investigated.
B.2 Scale-up and the sinusoidal hypothesis
The non-linear function of (2) records the memory behaviour of a quantum memristor element, when a sinusoidal input is used in input. Here, we want to both (i) scale-up the problem and (ii) compare the model $\mathcal{U}_{\rm sc+1}^{\rm BM}$ versus $\mathcal{U}_{\rm sc+1}^{\rm sin}$, bridging with SIREN classical networks. To address both questions, we display the learning curve for $11\mathcal{T}_{\rm ladder}$. For this experiment, we consider two topologies $\mathcal{U}_{\rm sc+1}^{\rm BM/sin}=\mathcal{U}_{\rm sc}^{\rm BM/sin}\otimes\mathbf{U}^{\rm BM/sin}_{10,0}$ and $\mathcal{W}_{\rm sc+1}^{\rm BM/sin}=\mathcal{U}_{\rm sc}\otimes\mathbf{U}^{\rm BM/sin}_{0,10}$.
Figure 8: Panel (a). Optimization history – the Meyer-Wallach entanglement maximization – for the $\mathcal{U}_{\rm sc+1}^{\rm BM}$ and $\mathcal{W}_{\rm sc+1}^{\rm BM}$ architecture with for a 11 qubits quantum network. Panel (b). The $\mathcal{U}_{\rm sc+1}^{\rm sin}$ and $\mathcal{W}_{\rm sc+1}^{\rm sin}$ architectures outcomes. The beam-splitter nonlinearity term brings a slightly better performance, but the sinsin term leads to a faster convergence. This may point out a better fitness of the sinesine non-linear term, for general cases, like the noisy one.
For 11 qubits, eq.2 and eq.3 are numerically equivalent. This means that sinusoidal non-linear functions are a valuable tool in our task of entanglement engineering, and the analogy with the SIREN is well-posed. The only appreciable difference between the two terms is the number of epochs (steps) needed to achieve convergence. The sinesine function offers a faster convergence; this can be the hallmark of a potential term with better fitness to our optimization problem. To conclude, we run the same analysis for 20 qubits either; this can be found in F.
Appendix C Initial condition dependency
Appendix D 5 noisy gates analysis
Figure 9: Histograms of 10 random topologies with amplitude noise on each wire after the $\mathrm{SWAP}$ gate that simulate the $\mathbf{U}^{\varkappa}$. Left histogram, $\mathcal{U}^{BM}_{\rm RN}$. right histogram, the $\mathcal{U}^{\theta}_{\rm RN}$ random topologies are compared. Also, for moderate noise, the better performances obtained from a non-linear function are evident with minimal sampling.
Even though the number of topologies here sampled is pretty low, we can immediately find one that can obtain entanglement values of $0.07$ and $0.31$ for the memristor type, and $0.09$ and $0.16$ for the sine non-linear element. The same outcomes are also confirmed for SC+1 in Fig.10.
Figure 10: Histogram for 10 different random topologies for a 5-qubit neural network. On the left, the $\mathcal{U}^{\rm sin}_{\rm SN}$ structure, and on the right, the $\mathcal{U}^{\rm\theta}_{\rm RN}$ model sharing same RN topologies. The sin model is pretty much equivalent to the beam-splitter, but always superior to the plain variational one.
We can notice again that the hypothesis that the sine non-linear components offer an evident advantage, and undoubtedly, for the global entanglement generation problem, this family of non-linearity bestows high efficiency. Compared to Fig.9 we can appreciate that both $\mathcal{U}_{\rm sc+1}^{\rm BM}$ and $\mathcal{U}_{\rm sc+1}^{\rm sin}$ are comparable.
Appendix E 10 qubits with amplitude damping noise
Figure 11: Panel (a). The 10 qubits $\mathcal{U}_{\rm sc+1}^{\rm BM}$ model for amplitude damping only with $\gamma=0.01$. The mild amount of noise does not hamper the global entanglement generation. Panel(b) the $\mathcal{U}_{\rm sc+1}^{\rm sin}$ ladder model. We can notice that the sinesine non-linear function provides faster convergence to the minima.
Appendix F 20 qubits experient
An open question is how the scaling in the number of qubits may affect the outcomes. In Fig.12 we compare the beam splitter non-linear function (upper panel) against the simple sinusoidal (lower panel). This simulation is provided for a very low amount of noise, with only amplitude damping after the two $\mathrm{SWAP}s$ for a $\gamma=10^{-3}$.
Figure 12: 20 qubit optimization. Upper panel, quantum network optimization with memristor-inspired reflectivity term. We can see that both topologies are equivalent. Lower panel, sinusoidal non-linear term application. We can check that the topology with the control term in the last wire (number zero) is bound to achieve a higher amount of entanglement.
For almost pure and pure states, the meristor-inspired term in (2) offers better performance in general. We have already detailed in Sec.II that this does not hold anymore for the realistic noisy case, where now the differential panorama is completely modified by the presence of local noise.

### v2

@@ Acknowledgements.
A.M.P. and R.L.F acknowledge support by MUR (Ministero dell’Università e della Ricerca) through the PNRR Project ICON-Q-Partenariato EstesoNQSTI - PE00000023 - Spoke 2 - CUP: J13C22000680006.
.1 Pure states random topologies analysis
Figure 6: Panels (a) (b). Histogram of the loss function achieved by a total of $10^{5}$ random $\mathcal{T}$ for 5 qubits. Panels (c) (d), histogram for 60 for 11 qubits. We compare the $\mathcal{U}_{\rm sc+1}^{\rm RN}$ (red bars) against $\mathcal{U}^{\theta}_{\rm RN}$ (yellow bars). The network loss function maximizes the MW values of the output states by minimizing Eq. 8. So, values close to zero in the loss function indicate higher values of global entanglement in the output pure state. We can clearly see that the presence of a nonlinearity as given by Eq. 16 response function, unlocks a greater optimal minima search.
As Fig.6 unfolds, the nonlinear components unlock better outcomes, with many more random QNN that can achieve values of MW above 0.8. This fact can be leveraged to prepare a training dataset for classical deep learning solutions meant to synthesize optimal topologies. We also run a quick analysis for a small noise scenario, where only amplitude damping is considered. Again, we confirm that the presence of nonlinear functions is necessary to achieve higher performances; the results are shown in Sec.B. Another natural question that arises is the dependency of the networks on initial conditions and parameter variation, i.e., the $T_{osc/int}$ that describes the physical gate, even though we are in a noiseless case. Results of the noiseless study can be found in A and show that for this scenario, no dependency is found.
The topologies of 0.8 threshold – As shown in panel (a), the MW value of 0.8 turns out as a threshold, or a specific boundary when a transition occurs. We analyze the structure of topologies below 0.8 and found that all of them have at least one wire set (A,B) with $A>B$ and that connect qubits that are not neighbors. This will play a critical role in Sec.  IV.2 when the presence of noise will be investigated.
.2 Scale-up and the sinusoidal hypothesis
The nonlinear function in Eq. (4) records the memory behaviour of a quantum memristor element, when a sinusoidal input is used in input. Here, we want to both (i) scale-up the problem and (ii) compare the model $\mathcal{U}_{\rm sc+1}^{\rm BM}$ versus $\mathcal{U}_{\rm sc+1}^{\rm sin}$, bridging with SIREN classical networks. To address both questions, we display the learning curve for $11\mathcal{T}_{\rm ladder}$. For this experiment, we consider two topologies $\mathcal{U}_{\rm sc+1}^{\rm BM/sin}=\mathcal{U}_{\rm sc}^{\rm BM/sin}\otimes\mathbf{U}^{\rm BM/sin}_{10,0}$ and $\mathcal{W}_{\rm sc+1}^{\rm BM/sin}=\mathcal{U}_{\rm sc}\otimes\mathbf{U}^{\rm BM/sin}_{0,10}$.
Figure 7: Panel (a). Optimization history – the Meyer-Wallach entanglement maximization – for the $\mathcal{U}_{\rm sc+1}^{\rm BM}$ and $\mathcal{W}_{\rm sc+1}^{\rm BM}$ architecture with for a 11 qubits quantum network. Panel (b). The $\mathcal{U}_{\rm sc+1}^{\rm sin}$ and $\mathcal{W}_{\rm sc+1}^{\rm sin}$ architectures outcomes. The beam-splitter nonlinearity term brings a slightly better performance, but the sinsin term leads to a faster convergence. This may point out a better fitness of the sinesine nonlinear term, for general cases, like the noisy one.
For 11 qubits, Eq. 4 and Eq. 5 are numerically equivalent. This means that sinusoidal nonlinear functions are a valuable tool in our task of entanglement engineering, and the analogy with the SIREN is well-posed. The only appreciable difference between the two terms is the number of epochs (steps) needed to achieve convergence. The sinesine function offers a faster convergence; this can be the hallmark of a potential term with better fitness to our optimization problem. To conclude, we run the same analysis for 20 qubits either; this can be found in D.
Appendix A Initial condition dependency
Appendix B 5 noisy gates analysis
Figure 8: Histograms of 10 random topologies with amplitude noise on each wire after the $\mathrm{SWAP}$ gate that simulate the $\mathbf{U}^{\varkappa}$. Left histogram, $\mathcal{U}^{BM}_{\rm RN}$. right histogram, the $\mathcal{U}^{\theta}_{\rm RN}$ random topologies are compared. Also, for moderate noise, the better performances obtained from a nonlinear function are evident with minimal sampling.
Even though the number of topologies here sampled is pretty low, we can immediately find one that can obtain entanglement values of $0.07$ and $0.31$ for the memristor type, and $0.09$ and $0.16$ for the sine nonlinear element. The same outcomes are also confirmed for SC+1 in Fig.9.
Figure 9: Histogram for 10 different random topologies for a 5-qubit neural network. On the left, the $\mathcal{U}^{\rm sin}_{\rm SN}$ structure, and on the right, the $\mathcal{U}^{\rm\theta}_{\rm RN}$ model sharing same RN topologies. The sin model is pretty much equivalent to the beam-splitter, but always superior to the linear one.
We can notice again that the hypothesis that the sine nonlinear components offer an evident advantage, and undoubtedly, for the global entanglement generation problem, this family of nonlinearity bestows high efficiency. Compared to Fig.8 we can appreciate that both $\mathcal{U}_{\rm sc+1}^{\rm BM}$ and $\mathcal{U}_{\rm sc+1}^{\rm sin}$ are comparable.
Appendix C 10 qubits with amplitude damping noise
Figure 10: Panel (a). The 10 qubits $\mathcal{U}_{\rm sc+1}^{\rm BM}$ model for amplitude damping only with $\gamma=0.01$. The mild amount of noise does not hamper the global entanglement generation. Panel(b) the $\mathcal{U}_{\rm sc+1}^{\rm sin}$ ladder model. We can notice that the sinesine nonlinear function provides faster convergence to the minima.
Appendix D 20 qubits experiment
An open question is how the scaling in the number of qubits may affect the outcomes. In Fig.11 we compare the beam splitter nonlinear function (upper panel) against the simple sinusoidal (lower panel). This simulation is provided for a very low amount of noise, with only amplitude damping after the two $\mathrm{SWAP}s$ for a $\gamma=10^{-3}$.
Figure 11: 20 qubit optimization. Upper panel, quantum network optimization with memristor-inspired reflectivity term. We can see that both topologies are equivalent. Lower panel, sinusoidal nonlinear term application. We can check that the topology with the control term in the last wire (number zero) is bound to achieve a higher amount of entanglement.
For almost pure and pure states, the memristor-inspired term in Eq. (4) offers better performance in general. We have already detailed in Sec.III that this does not hold anymore for the realistic noisy case, where now the differential panorama is completely modified by the presence of local noise.
@@ Appendix A Initial condition dependency
In recent years, a revival of photonic applications to extracting quantum state properties (or features) is showing a resurgence, with several works on photonic platforms [ExperimentalPropertiesPhotonicPlatform, ReviewReservoir-Innocenti2023, Reservoir-Spagnolo2022].
Appendix E The beam splitter reflectivity
@@
Appendix F Structure of the Response Function
@@
$\omega=\frac{2\pi}{T_{\mathrm{osc}}},$
@@
$\frac{T_{\mathrm{osc}}}{T_{\mathrm{int}}\,4\pi}=\frac{1}{2\omega T_{\mathrm{int}}},$
@@
In systems exhibiting memory-dependent (or “memristive”) behavior, the present response does not depend solely on the instantaneous value of the driving field, but also on its history within a specific temporal window. Eq. (15) captures exactly this effect: $R(t)$ measures the deviation introduced by integrating the oscillatory signal over an interval $T_{\mathrm{int}}$, effectively encoding a temporal “lag” between the free oscillation and its memory-processed counterpart.
@@
This formulation offers a compact yet powerful tool to quantify the interplay between coherence, periodic driving, and memory depth. In the context of photonic or quantum memristive systems, the response function Eq. (15) is particularly useful for characterizing the temporal retention of coherence and for distinguishing between purely dissipative and genuinely history-dependent dynamics.